• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name J. Butler (Tech)

    Log entry time 11:27:04 on March 14, 2014

    Entry number 386760

    keyword=Pre-Beam Checklist

    Checklist complete

    The target, spectrometer entrance, spectrometer exit and VDC covers are all installed.

    Angle limits for the left HRS are "15 - 15 degrees".

    Angle limits for the right HRS are "33 - 45 degrees".

    The left and right snout collimator has been removed.

    Both HRS spectrometers are down for troubleshooting of power supplies (inquiries contact Jack Segal)

    Lead flow may be unstable while troubleshooting is in progress.

    Tech on call is Jessie Butler.



    A copy of this log entry has been emailed to: folts@jlab.org, jbutler@jlab.org